OpenStudy (anonymous):

What is the real and imaginary parts of the complex number -5+6i?

OpenStudy (anonymous):

A complex number is written in the form of a+bi, where a is the real part and b is the imaginary part.

OpenStudy (anonymous):

Can you also help me with multiplying complex numbers? I am not good with it. For example, i(2+i) equals what? And (1+2i)-(1+5i) equals what? And then there's simplifying them. Grrrrrrr.

OpenStudy (anonymous):

To multiply (2+3i) with (5-2i) you do: \[(2+3i)*(5-2i) = 2*5 - 2*2i + 3i*5 - 2i*3i\] Then you group and replace i^2 with -1. The subtraction is done in the same way with addition. In general, you treat i as a constant. The difference is that it's square equals -1.

OpenStudy (anonymous):

So i equals to a number that's actually not a number? So, how do you simplify complex numbers then? \[\frac{ 5+2i }{ 6+i }\]

OpenStudy (anonymous):

i is the imaginary unit. I told you to treat it as a constant, so you will multiply (2+3i) with (5-2i) as if you had (2+3a) with (5-2a), where a is a constant.

OpenStudy (anonymous):

About the division \[\frac{ 5+2i }{ 6+i }\] you want to get rid of the 6+i. You want to make it real. So what is the result if you multiply (6+i) with (6-i)?

OpenStudy (anonymous):

Oh, oh! I get it now! Oh my goodness, thank you! Yes!

OpenStudy (anonymous):

Did you find the result of the multiplication (6+i) with (6-i) ?

OpenStudy (anonymous):

Is it 37?

OpenStudy (anonymous):

That's right! Bravo! Can you see now what happens to \[\frac{ 5+2i }{ 6+i }\] when you multiply both numerator and denominator with (6-i)?